A jet plane lands with a velocity of +105 m/s and can accelerate at a maximum rate of -5.13 m/s2 as it comes to rest. From the instant it touches the runway, what is the minimum time needed before it can come to rest?
Answer #1

Vo= initial speed, Vf= final speed
Vf = Vi + at
0 = 105 -5.13t
5.13t= 110
t = 110/5.13

t = 105/5.13 = 20.4678 s to stop